OpenStudy (anonymous):

e first correct answer will get a medal!!!! which is a solution of the equation? 2x-7=5 a.)(2,-1) b.)7 c.)(3,-1) d.)1,2)

OpenStudy (anonymous):

2x - 7 = 5 2x = 5 + 7 2x = 12 (divide both sides by 2 to get x) 2x/2 = 12/2 x = 6

OpenStudy (anonymous):

nope I need the solution

OpenStudy (the_fizicx99):

@mangorox you're not solving the equation your plugging in the values to see which will = 5

OpenStudy (the_fizicx99):

There's a piece missing from either your options or the equation itself

OpenStudy (anonymous):

THis question makes no sense